OpenStudy (anonymous):

\[\sum _{k=-\infty }^{\infty } \theta (n-k)\theta (-k-1)a^k\]

OpenStudy (anonymous):

\[\theta\] is step function

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

I can show you the work , it would be great if you can explain it

OpenStudy (anonymous):

I get the n<=-1 part but not the n>-1

OpenStudy (anonymous):

this is convultion

OpenStudy (phi):

I know u(-k-1) is 1 for k= -inf to -1 u(x) is 0 for x<0, and 1 for x≥0 u(n-k) is 1 for k= -inf to +n

OpenStudy (anonymous):

use convolution property

OpenStudy (anonymous):

you can break up the sequence from -infinity to 0 and from 0 to infinity for 0 to infinity we know what are the results of a step function

OpenStudy (phi):

when you multiply the 2 step functions the one that "steps down" (because we are doing u(-k) ) first sets the upper limit n wins for n< -1, and the other wins for n> -1

OpenStudy (phi):